OpenStudy (anonymous):

Help!!!!! 2 + 11b= 8b + 15

OpenStudy (anonymous):

Group like terms first. meaning subtract 2 from both sides and sub tract 8b from both sides.

OpenStudy (anonymous):

Then solve for b

OpenStudy (anonymous):

i got 13 and 3

OpenStudy (anonymous):

if you do the same thing to both sides of an equation, the equation essentially doesnt change it just was manipulated a bit.

OpenStudy (anonymous):

3b = 13, next step is to divide both sides by 3 to get b by itself

OpenStudy (anonymous):

11b - 8b = 15 - 2 3b = 13 b = 13 / 3 b =4.33

OpenStudy (anonymous):

3/3 = 1 so your left with b = ?

OpenStudy (anonymous):

so i divide 3 by 13 which is a fraction 4 1/3

OpenStudy (anonymous):

13 by 3, but yes you got the correct answer

OpenStudy (anonymous):

thanks so much everyone (: These problems are so confusing!!!!

OpenStudy (anonymous):

thats equal. . . 13 / 3 = 4 1/3 thats improper and proper fraction . try to review primary topics i think grade 1?

OpenStudy (anonymous):

@depads don't put people down. Sometimes things don't click and we need to help them through it not just tell them the answer.